OpenStudy (anonymous):

Help please? =) I always medal/fan those that help =) Simplify. ((24u^6x^4 - 18u^4x^4 + 12u^5x^3)) / ((3u^5x^3))

OpenStudy (anonymous):

I got this " (8 u^2 x+4 u-6 x)/u " but it says I need to simplify it further and I need help =o

OpenStudy (anonymous):

@YOLO_KING

OpenStudy (anonymous):

I got you!!!!

OpenStudy (anonymous):

I can help you. but I can only give an example. Is that okay with you?

OpenStudy (anonymous):

EXAMPLE: 1) Find the greatest common factor of the two expression: Simplify 20uy^8z^6 and 28u^7y^4 --- gcf for 20 and 28 is 4 gcf for u and u^7 is u gcf for y^8 and y^4 is y^4 gcf for z^6 and no z is no z --------------------------------- GCF = 4uy^4 ======================================================================== 2) Find the least common multiple of the two expressions: Simplify 3t^6x^8y^5 and 12x^3y^6 lcm of 3 and 12 is 12 lcm of t^6 and no t is t^6 lcm of x^8 and x^3 is x^8 lcm of y^5 and y^6 is y^6 --- LCM = 12t^6*x^8*y^6 ======================================================== 3) Factor Completely: 27z^3+125 No common factor ------------------------------- 4) Factor: 9t+wt+9y+yw --- t(9+w) + y(9+w) --- = (9+w)(t+y) ------------------------------- 5) Factor: -8t+xt-24x+3x^2 --- = -(8t -xt + 24x - 3x^2) --- = -(t(8-x) + 3x(8-x)) --- = -(8-x)(t+3x)

OpenStudy (anonymous):

\[\frac{ 3u^4x^3(8u^2x-6x+4u) }{ 3u^5x^3 }\] I factor out so now you can cross them out

OpenStudy (anonymous):

you should get 8ux-(6x/u)+4

OpenStudy (anonymous):

Good job^

OpenStudy (anonymous):

Thank you both so much!!! =D